If the satellite S2 at a distance 6850000 m from the center of the earth, determined the orbital period of the satellite.

[Mass of earth, M = 5.97x 〖10〗^24kg, G=6.67x〖10〗^(-11) 〖Nm〗^2 〖kg〗^(-2)


1
Expert's answer
2021-12-20T10:27:11-0500

Explanations & Calculations


  • For this, you need to consider the equation used for the gravitation between two bodies.

"\qquad\qquad
\begin{aligned}
\small F&=\small ma\cdots[\text{to the satellite towards the center of its orbit}]\\
\small G\frac{M_em_s}{r^2}&=\small m_sr\omega^2\\
\small G\frac{M_e}{r^2}&=\small r\omega^2\\
\small G\frac{M_e}{r^2}&=\small r(\frac{2\pi}{T})^2\\
\small T^2&=\small4\pi^2.r.\frac{r^2}{GM_e}\\
\small T^2&=\small 4\pi^2.\frac{r^2}{GM_e} \\
\small T&=\small \sqrt{4\pi^2.\frac{r^3}{GM_e}}\\
&=\small\sqrt{4\pi^2.\frac{(6.85\times10^6\,m)^3}{(6.67\times10^{-11}Nm^2kg^{-2})(5.97\times10^{27}\,kg)}}\\
&=\small \sqrt{31866.27\times\frac{m^3}{kgms^{-2}.m^2.kg^{-2}.kg}}\\
&=\small \sqrt{31866.27\,s^2}\\
&=\small 178.51\,s
\end{aligned}"

  • You need to subject the quantity you need to get an answer for, once you know the equation and substitute the values and get the answer.
  • I have provided a step by step guide to the final answer.
  • Newtons need to be written in SI units to figure out the final unit.



  • I believe you can substitute the given values into the equation and get the answer if that was the case no point in posting this question.
  • Next time you post one, do mention where you really lag in getting the answer, is it the equation, any calculation or need to recheck the answer you already got.